# Leaf schemes and Hodge loci [^1] February 25, 2025

### v1

@@ 1.3 Bad primes
The number $N\in{\mathfrak{R}}$ in 1.1 is the main responsible for the finite number of exceptional primes in the Grothendieck-Katz conjecture. In this section we describe these primes.  [^4] For examples of bad primes of Lamé equations see Section 2.
@@
If ${\mathfrak{R}}=\mathbb{Z}$ then this proposition follows from the unique factorization. The bad primes are those dividing $N$. Let us now consider a number field ${\mathfrak{k}}$ and its ring of integers ${\mathfrak{R}}={\cal O}_{\mathfrak{k}}$. By unique factorization theorem for ideals in ${\cal O}_{\mathfrak{k}}$, we know that there are only a finite number of prime ideals in ${\cal O}_{\mathfrak{k}}$ containing $N{\cal O}_{\mathfrak{k}}$. Let $p_{1},p_{2},\ldots,p_{s}\in\mathbb{Z}$ be the list of characteristics of the corresponding residue fields. Take an arbitrary $m\in\mathbb{Z}$ coprime with $p_{i}$’s and such that $m|N^{d}M$ in ${\mathfrak{R}}$. By construction $m$ and $N^{d}$ are coprime, that is, there is no prime ideal containing both $m{\cal O}_{\mathfrak{k}}$ and $N^{d}{\cal O}_{\mathfrak{k}}$. Therefore, $m{\cal O}_{\mathfrak{k}}+N^{d}{\cal O}_{\mathfrak{k}}={\cal O}_{\mathfrak{k}}$. In particular, there is $a,b\in{\cal O}_{\mathfrak{k}}$ such that $ma+N^{d}b=1$. Multiplying this equality with $M$ we conclude that $M\in m{\cal O}_{\mathfrak{k}}$. ∎
@@ 1.4 Algebraicity of a single solution
[…]
@@
The proof is the same as the proof of 1.1. In the last step (3), instead of $Y$ we use $y$ itself and evaluate it in $z=z_{0}$. We have also $\Delta(z_{0})\in{\mathfrak{R}}$ which is absorbed by $N$, that is, instead of ${\mathfrak{R}}[\frac{1}{N\Delta(z_{0})}]$ we have written ${\mathfrak{R}}[\frac{1}{N}]$. It is already evident what is the analog of 1.1 in this framework: If for a solution $y(z)$ of $\frac{\partial y}{\partial z}={\sf A}y$ with $y(z_{0})=y_{0}$ we have (5) then $y(z)$ must be algebraic.  [^5] Daniel Litt and Joshua Lam kindly reminded me that this conjecture is related to a conjecture of Y. André and G. Christol in [And04, Remarque 5.3.2]. In [BCR24, Conjecture 1.12] this is actually attributed to them. The analog of Grothendieck-Katz conjecture is: if for all but a finite number of primes $p$, ${\sf A}_{p}(z_{0})y_{0}$ vanishes modulo $p$ then $y(z)$ is algebraic.  [^6] This statement is false. See 3.2.
@@ 1.5 Integral functions
(6) $(y_{1}y_{1})^{[n]}$ $=$ $\sum_{i=0}^{n}y_{1}^{[i]}y_{2}^{[n-i]},$
(7) $(y^{-1})^{[n]}$ $=$ $-y^{-1}\sum_{i=0}^{n-1}y^{[n-i]}(y^{-1})^{[i]},$
(8) $(y^{m})^{[n]}$ $=$ $\sum_{i_{1}+i_{2}+\cdots+i_{m}=n}y^{[i_{1}]}y^{[i_{2}]}\cdots y^{[i_{m}]},$
(9) $(y^{[n]})^{[m]}$ $=$ $\binom{n+m}{n}y^{[n+m]},$
(10) $(z^{a})^{[n]}$ $=$ $[a]_{n}z^{a-n},\ \ \hbox{In particular }(z^{-1})^{[n]}=(-1)^{n}z^{-n-1}.$
@@
(11) $\mathbb{Z}[x]\to\mathbb{Z}[x],\ \ P\mapsto P^{[n]},$
@@
The operator (11) is $\mathbb{Z}$-linear, and so in order to show that $\mathbb{Z}[x]$ is closed under (11), we need to show it for a monomial in $x$. From (6) it follows that in order to verify that $\mathbb{Z}[x]$ is an integral algebra it is enough to prove that $x_{i}^{[n]}\in\mathbb{Z}[x]$ for all $n\in\mathbb{N}_{0}$ and for a set of generators $x=(x_{1},x_{2},\ldots,x_{N})$ of $\mathbb{Z}[x]$.
@@
If $\mathbb{Z}[x]$ and $\mathbb{Z}[\tilde{x}]$ are integral $\mathbb{Z}$-algebra, then by 1.2, $\mathbb{Z}[x,\tilde{x}]$ is also integral. This implies that if $y$ and $\tilde{y}$ are two integral functions with the associated $\mathbb{Z}$-algebras $\mathbb{Z}[x]$ and $\mathbb{Z}[\tilde{x}]$ then $y+\tilde{y},\ y\tilde{y}$ are also integral with the associated $\mathbb{Z}$-algebra $\mathbb{Z}[x,\tilde{x}]$. For an integral function $y\in\mathbb{Z}[x]$, $\mathbb{Z}[y^{-1},x]$ is also an integral algebra. This follows from (7) and 1.2. ∎
@@
(12) $y^{[k]}\in\mathbb{Z}[x]:=\mathbb{Z}\left[\left(\frac{\partial P}{\partial y}\right)^{-1},\ z,{\rm coef}(P),\ y^{[i]},\ \ \ i=0,1,\ldots,d\right],\ \ k\geq d+1,$
where ${\rm coef}(P)$ is the list of coefficients of $P$ (they might be transcendendtal numbers). By 1.2 we need to show $P^{[n]}\in\mathbb{Z}[x]$ for generators of $\mathbb{Z}[x]$. For $P=\left(\frac{\partial P}{\partial y}\right)^{-1}$ this follow from (7). For $P=y^{[n]}$ this follows from (9) and (12). ∎
@@
It is clear from (12) the the length of an algebraic function $y$ is less that or equal
[…]
@@
[…]
@@ 1.6 An attempt
The author’s impression is that 1.3 which is a consequence of Grothendieck-Katz conjecture might be false. In order to transfer this feeling to the reader, in this section we attempt to translate the vanishing of $p$-curvatures into a characteristic zero statement. Let ${\mathfrak{R}}$ be the $\mathbb{Z}$-algebra generated by the coefficients of ${\sf A}$ and ˇ$\check{\mathfrak{R}}={\mathfrak{R}}[z_{0},\frac{1}{\Delta(z_{0})},y_{0}]$ for $y_{0}\in\mathbb{C}^{\sf n}$, $z_{0}\in\mathbb{C}$ which is not a pole of ${\sf A}$.
Theorem 1.4.
If the $p$-curvature of $y^{\prime}={\sf A}y$ is zero then we have a formal power series $\check{y}\in\check{\mathfrak{R}}[[z-z_{0}]]$ with $\check{y}=y_{0}$ such that $\check{y}^{\prime}={\sf A}\check{y}$ in ˇ$\check{\mathfrak{R}}/p\check{\mathfrak{R}}$.
@@
The proof is a simplification and adaptation of [Ses60, Theorem 2] into our context. Let ${\mathfrak{R}}_{p}:={\mathfrak{R}}/p{\mathfrak{R}}$. We consider the larger ring ${\mathfrak{R}}_{p}[z,y]$ and the derivation
@@
If not, there is a $f\in{\mathfrak{R}}_{p}[z,y]$ such that $f\not\in{\rm ker}_{p}({\sf v})[z]$. Since ${\sf v}^{p}f=0$, for some $11\leq e\leq p-$ 1 we have $g={\sf v}^{e}f\not\in{\rm ker}_{p}({\sf v})[z]$ but ${\sf v}g\in{\rm ker}_{p}({\sf v})[z]$. Let us write
@@
Let us take $y_{0}\in\mathbb{C}^{\sf n}$ and $z_{0}\in\mathbb{C}$ which is not a pole of ${\sf A}$. From now on we use the larger ring ˇ$\check{\mathfrak{R}}={\mathfrak{R}}[y_{0},z_{0},\frac{1}{\Delta(z_{0})}]$ and the equalities are in ˇ$\check{\mathfrak{R}}_{p}:=\check{\mathfrak{R}}/p\check{\mathfrak{R}}$. We write
@@
This equality implies that the algebraic variety $a_{0}(z,y)=0$ defined over ˇ$\check{\mathfrak{R}}_{p}$ contains the point $(y,z)=(y_{0},z_{0})$. Actually, it is an algebraic in $\mathbb{A}^{n+1}_{\check{\mathfrak{R}}_{p}}$ and by definition of ${\rm ker}_{p}({\sf v})$, it is tangent to the vector field ${\sf v}$. We get the map
@@
whose derivative at $(z_{0},y_{0})$ has the determinant $1$. By inverse function theorem over a ring, we have power series $\check{y}\in\check{\mathfrak{R}}_{p}[[z-z_{0},w]]$ such that $a_{0}(z,\check{y})=w$. We redefine $\check{y}$ to be $\check{y}$ restricted to $w=0$ and get $\check{y}\in\check{\mathfrak{R}}_{p}[[z-z_{0}]]$. Therefore, $a_{0}(z,\check{y})=0$. This together with ${\sf v}(a_{0})=0$, imply that
@@
We take a representative for the coefficients of $y(p):=\check{y}$ in $\check{\mathfrak{R}}$ and use $y(p)$ to emphasize that it depends on $p$. Therefore, $y(p)\in\check{\mathfrak{R}}[[z-z_{0}]]^{\sf n}$ and $\frac{\partial y(p)}{\partial z}\equiv_{p}{\sf A}(z)y(p)$. The main difficulty in proving that the entries of $y$ are integral functions is to glue the data of all $y(p)$ for different $p$ together. By 1.4 we know that over $\check{\mathfrak{R}}\otimes_{\mathbb{Z}}\mathbb{Q}$ we have a unique solution $y$ with with $y(z_{0})=y_{0}$ (here we must assume that $\Delta(z_{0})$ is invertible in $\check{\mathfrak{R}}$). We might claim that $y$ has coefficients in $\check{\mathfrak{R}}$ and $y\equiv_{p}y(p)$ and try to prove this by induction on $n$ for the coefficient of $(z-z_{0})^{n}$. This is trivially true for coefficients of $(z-z_{0})^{i},\ \ i=0,1$. Assume that it is true for all $m<n$. The coefficient of $(z-z_{0})^{n}$ in $y$ is computed by
(14) $ny_{n}=\sum_{i+j=n-1}{\sf A}_{i}y_{j}.$
@@
We have also $ny(p)_{n}\equiv_{p}\sum_{i+j=n-1}{\sf A}_{i}y(p)_{j}$. For $n\leq 2p-1$, the right hand side of (14) is divisible by $p$ and so $y_{n}$ can be computed in $\check{\mathfrak{R}}$. However, for $n=p^{2}$, we observe that $y_{n}$ might have $p$ in its denominator and the argument breaks. One might try to give a recursive formula for the order of $p$ in the denominator of $y_{n}$.

### v2

@@ 1.3 Bad primes
The number $N\in{\mathfrak{R}}$ in 1.1 is the main responsible for the finite number of exceptional primes in the Grothendieck-Katz conjecture. In this section we describe these primes.  [^2] For examples of bad primes of Lamé equations see Chapter 2.
@@
If $\mathfrak{R}$$\mathbb{Z}$ then this proposition follows from the unique factorization. The bad primes are those dividing $N$. Let us now consider a number field $\mathfrak{k}$} and its ring of integers ${\mathfrak{R}}={\mathcal{O}}_{\mathfrak{k}}$. By unique factorization theorem for ideals in ${\mathcal{O}}_{\mathfrak{k}}$, we know that there are only a finite number of prime ideals in ${\mathcal{O}}_{\mathfrak{k}}$ containing $N{\mathcal{O}}_{\mathfrak{k}}$. Let $p_{1},p_{2},\ldots,p_{s}\in\mathbb{Z}$ be the list of characteristics of the corresponding residue fields. Take an arbitrary $m\in\mathbb{Z}$ coprime with $p_{i}$’s and such that $m|N^{d}M$ in $\mathfrak{R}$}. By construction $m$ and $N^{d}$ are coprime, that is, there is no prime ideal containing both $m{\mathcal{O}}_{\mathfrak{k}}$ and $N^{d}{\mathcal{O}}_{\mathfrak{k}}$. Therefore, $m{\mathcal{O}}_{\mathfrak{k}}+N^{d}{\mathcal{O}}_{\mathfrak{k}}={\mathcal{O}}_{\mathfrak{k}}$. In particular, there is $a,b\in{\mathcal{O}}_{\mathfrak{k}}$ such that $ma+N^{d}b=1$. Multiplying this equality with $M$ we conclude that $M\in m{\mathcal{O}}_{\mathfrak{k}}$. ∎
@@ 1.4 Algebraicity of a single solution
(1.5)  $\hbox{ entries of }\frac{N^{m}{\sf A}_{m}(z_{0})y_{0}}{m!}\ \in\ {\mathfrak{R}},\ \ \forall m\in\mathbb{N}.$
@@
The proof is the same as the proof of 1.1. In the last step (1.3), instead of $Y$ we use $y$ itself and evaluate it in $z=z_{0}$. We have also $\Delta(z_{0})\in{\mathfrak{R}}$ which is absorbed by $N$, that is, instead of ${\mathfrak{R}}[\frac{1}{N\Delta(z_{0})}]$ we have written ${\mathfrak{R}}[\frac{1}{N}]$. It is already evident what is the analog of 1.1 in this framework: If for a solution $y(z)$ of $\frac{\partial y}{\partial z}={\sf A}y$ with $y(z_{0})=y_{0}$ we have (1.5) then $y(z)$ must be algebraic.  [^3] Daniel Litt and Joshua Lam kindly reminded me that this conjecture is related to a conjecture of Y. André and G. Christol in [Andre2004, Remarque 5.3.2]. In [BostanCarusoRoques2024, Conjecture 1.12] this is actually attributed to them. The analog of Grothendieck-Katz conjecture is: if for all but a finite number of primes $p$, ${\sf A}_{p}(z_{0})y_{0}$ vanishes modulo $p$ then $y(z)$ is algebraic.  [^4] This statement is false. See 3.2.
@@ 1.5 Integral functions
(1.6) $(y_{1}y_{1})^{[n]}$ $=$ $\sum_{i=0}^{n}y_{1}^{[i]}y_{2}^{[n-i]},$
(1.7) $(y^{-1})^{[n]}$ $=$ $-y^{-1}\sum_{i=0}^{n-1}y^{[n-i]}(y^{-1})^{[i]},$
(1.8) $(y^{m})^{[n]}$ $=$ $\sum_{i_{1}+i_{2}+\cdots+i_{m}=n}y^{[i_{1}]}y^{[i_{2}]}\cdots y^{[i_{m}]},$
(1.9) $(y^{[n]})^{[m]}$ $=$ $\binom{n+m}{n}y^{[n+m]},$
(1.10) $(z^{a})^{[n]}$ $=$ $[a]_{n}z^{a-n},\ \ \hbox{In particular }(z^{-1})^{[n]}=(-1)^{n}z^{-n-1}.$
@@
(1.11) $\mathbb{Z}[x]\to\mathbb{Z}[x],\ \ P\mapsto P^{[n]},$
@@
The operator (1.11) is $\mathbb{Z}$-linear, and so in order to show that $\mathbb{Z}[x]$ is closed under (1.11), we need to show it for a monomial in $x$. From (1.6) it follows that in order to verify that $\mathbb{Z}[x]$ is an integral algebra it is enough to prove that $x_{i}^{[n]}\in\mathbb{Z}[x]$ for all $n\in\mathbb{N}_{0}$ and for a set of generators $x=(x_{1},x_{2},\ldots,x_{N})$ of $\mathbb{Z}[x]$.
@@
If $\mathbb{Z}[x]$ and $\mathbb{Z}[\tilde{x}$] are integral $\mathbb{Z}$-algebra, then by 1.2, $\mathbb{Z}[x,\tilde{x}$] is also integral. This implies that if $y$ and $\tilde{y}$ are two integral functions with the associated $\mathbb{Z}$-algebras $\mathbb{Z}[x]$ and $\mathbb{Z}[\tilde{x}$] then $y+\tilde{y},\ y\tilde{y}$ are also integral with the associated $\mathbb{Z}$-algebra $\mathbb{Z}[x,\tilde{x}$]. For an integral function $y\in\mathbb{Z}[x]$, $\mathbb{Z}[y^{-1},x]$ is also an integral algebra. This follows from (1.7) and 1.2. ∎
@@
(1.12) $y^{[k]}\in\mathbb{Z}[x]:=\mathbb{Z}\left[\left(\frac{\partial P}{\partial y}\right)^{-1},\ z,{\rm coef}(P),\ y^{[i]},\ \ \ i=0,1,\ldots,d\right],\ \ k\geq d+1,$
where ${\rm coef}(P)$ is the list of coefficients of $P$ (they might be transcendendtal numbers). By 1.2 we need to show $P^{[n]}\in\mathbb{Z}[x]$ for generators of $\mathbb{Z}[x]$. For $P=\left(\frac{\partial P}{\partial y}\right)^{-1}$ this follow from (1.7). For $P=y^{[n]}$ this follows from (1.9) and (1.12). ∎
@@
It is clear from (1.12) the the length of an algebraic function $y$ is less that or equal
$N:=d+3+\hbox{number of monomials in $P$ with non-zero coefficients}-1.$
@@
(1.13)  $\hbox{ entries of }\frac{y^{(m)}}{m!}\ \in\ {\mathfrak{R}}[z,\frac{1}{N\Delta},y_{1},y_{2},\ldots,y_{n}],\ \ \forall m\in\mathbb{N}.$
@@
Remark 1.5.
Let us consider a differential equations $y^{\prime}={\sf A}y$ satisfying the hypothesis of 1.1. It is a straightforward computation to show that
$U=\sum_{k=0}^{\infty}\frac{(-1)^{k}{\sf A}_{k}(z)}{k!}(z-z_{0})^{k},$
with $U(0)=$ the identity matrix, satisfies the differential equation $U^{\prime}=-U{\sf A}$, and hence, $Y:=U^{-1}$ is a fundamental matrix of $Y^{\prime}={\sf A}Y$, see [BostanCarusoRoques2024, Remark 3.20] . We can consider $U$ and $U^{-1}$ as formal power series in $(z-z_{0})$ and the coefficients are in the ring ${\mathfrak{R}}[z_{0},\frac{1}{N\Delta(z_{0})}]$ . Note that we have also
$U=\sum_{k=0}^{\infty}\frac{{\sf A}_{k}(z_{0})}{k!}(z-z_{0})^{k},$
which follows again by the hypothesis of 1.1 and the formula for Taylor series.
@@ 1.6 An attempt
The author’s impression is that 1.3 which is a consequence of Grothendieck-Katz conjecture might be false. In order to transfer this feeling to the reader, in this section, we attempt to translate the vanishing of $p$-curvatures into the $p$-integrality of its solutions. T Let $\mathfrak{R}$} be the $\mathbb{Z}$-algebra generated by the coefficients of $\sf A$} and $\check{\mathfrak{R}}={\mathfrak{R}}[z_{0},\frac{1}{\Delta(z_{0})},y_{0}]$ for $y_{0}\in\mathbb{C}^{\sf n}$, $z_{0}\in\mathbb{C}$ which is not a pole of $\sf A$}.
Proposition 1.8.
If the $p$-curvature of $y^{\prime}={\sf A}y$ is zero then we have a formal power series $\check{y}\in\check{\mathfrak{R}}[[z-z_{0}]]$ with $\check{y}=y_{0}$ such that $\check{y}^{\prime}={\sf A}\check{y}$ in $\check{\mathfrak{R}}/p\check{\mathfrak{R}}$.
@@
We give two proofs of this statement. The first proof is well-known, see [BostanCarusoRoques2024, Remark 3.20], and it is less intuitive. It is a mere computation to see that if ${\sf A}_{p}=0$ in $\check{\mathfrak{R}}/p\check{\mathfrak{R}}$ then the sum $U=\sum_{k=0}^{p-1}\frac{(-1)^{k}}{k!}(z-z_{0})^{k}{\sf A}_{k}$ with $U(0)=$ identity matrix, satisfies the differential equation $U^{\prime}=-U{\sf A}$, and hence, $U^{{\mathsf{t}}{\mathsf{r}}}$ is the solution of the dual differential equation $y^{\prime}=-{\sf A}^{{{\mathsf{t}}{\mathsf{r}}}}y$. It turns out that $U^{-1}$ is a fundamental matrix of $y^{\prime}={\sf A}y$.
The second proof is more natural and it is a simplification and adaptation of [Seshadri1960, Theorem 2] into our context. Let ${\mathfrak{R}}_{p}:={\mathfrak{R}}/p{\mathfrak{R}}$. We consider the larger ring ${\mathfrak{R}}_{p}[z,y]$ and the derivation
@@
If not, there is a $f\in{\mathfrak{R}}_{p}[z,y]$ such that $f\not\in{\rm ker}_{p}({\sf v})[z]$. Since ${\sf v}^{p}f=0$, for some $1\leq e\leq p-1$ we have $g={\sf v}^{e}f\not\in{\rm ker}_{p}({\sf v})[z]$ but ${\sf v}g\in{\rm ker}_{p}({\sf v})[z]$. Let us write
@@
Let us take $y_{0}\in\mathbb{C}^{\sf n}$ and $z_{0}\in\mathbb{C}$ which is not a pole of $\sf A$}. From now on we use the larger ring $\check{\mathfrak{R}}={\mathfrak{R}}[y_{0},z_{0},\frac{1}{\Delta(z_{0})}]$ and the equalities are in $\check{\mathfrak{R}}_{p}:=\check{\mathfrak{R}}/p\check{\mathfrak{R}}$. We write
@@
This equality implies that the algebraic variety $a_{0}(z,y)=0$ in $\mathbb{A}^{n+1}_{\check{\mathfrak{R}}_{p}}$ contains the point $(y,z)=(y_{0},z_{0})$. By definition of ${\rm ker}_{p}({\sf v})$, it is tangent to the vector field $\sf v$}. We get the map
@@
whose derivative at $(z_{0},y_{0})$ has the determinant $1$. By inverse function theorem over a ring, see 1.6, we have power series $\check{y}\in\check{\mathfrak{R}}_{p}[[z-z_{0},w]]$ such that $a_{0}(z,\check{y})=w$. We redefine $\check{y}$ to be $\check{y}$ restricted to $w=0$ and get $\check{y}\in\check{\mathfrak{R}}_{p}[[z-z_{0}]]$. Therefore, $a_{0}(z,\check{y})=0$. This together with ${\sf v}(a_{0})=0$, imply that
@@
The inverse function theorem and implicit function theorem are true over an arbitrary ring $\mathfrak{R}$} as follows. Let $A=\mathbb{A}_{{\mathfrak{R}}}^{n}={\rm Spec}({\mathfrak{R}}[z_{1},z_{2},\ldots,z_{n}])$ and ${t}=(t_{1},t_{2},\ldots,t_{n}),\ {t}_{i}\in{\mathfrak{R}}$ be an $\mathfrak{R}$}-valued point of $A$. Let also ${\mathcal{O}}_{A^{\rm for},{t}}$ (resp. ${\mathcal{O}}_{A^{\rm hol},{t}}$) be the ring of formal (resp. convergent for some embedding $\mathfrak{R}$$\subset\mathbb{C}$) power series $\sum_{i\in\mathbb{N}_{0}^{n}}a_{i}(z-t)^{i}$.
  1. 1.
    Let $F=(F_{1},F_{2},\ldots,F_{n})$ with $F_{i}\in{\mathcal{O}}_{A^{\rm hol},{t}}$ (resp. ${\mathcal{O}}_{A^{\rm for},{t}}$). If $\det[\frac{\partial F_{i}}{\partial z_{j}}(0)]$ is invertible in $\mathfrak{R}$} then there is $G=(G_{1},G_{2},\ldots,G_{n})$, with $G_{i}\in{\mathcal{O}}_{A^{\rm hol},{t}}$ (resp. ${\mathcal{O}}_{A^{\rm for},{t}}$) such that $F\circ G=G\circ F=(z_{1},z_{2},\ldots,z_{n})$.
  2. 2.
    Let $F=(F_{1},F_{2},\ldots,F_{m})$ with $F_{i}\in{\mathcal{O}}_{A^{\rm hol},{t}}$ (resp. ${\mathcal{O}}_{A^{\rm for},{t}}$). If the rows of $\frac{\partial F_{i}}{\partial z_{j}}(0)]$ can be completed to a basis of ${\mathfrak{R}}^{n}$ then there is $G=(G_{1},G_{2},\ldots,G_{n-m})$, with $G_{i}\in{\mathcal{O}}_{B^{\rm hol},0}$ (resp. ${\mathcal{O}}_{B^{\rm for},{t}}$) and $B=\mathbb{A}_{{\mathfrak{R}}}^{n-m}={\rm Spec}({\mathfrak{R}}[z_{1},z_{2},\ldots,z_{n-m}])$ such that $\frac{\partial G_{i}}{\partial z_{j}}(0)]$ is invertible in $\mathfrak{R}$} and
    $F(\check{z},G(\check{z}))=0,$
    where $\check{z}=(z_{1},z_{2},\ldots,z_{n-m})$.
Definition 1.2.
Let $\mathfrak{R}$} be a finitely generated subring of $\mathbb{C}$ such that a primes number $p$ is not invertible in $\mathfrak{R}$}. We say that a formal power series is $p$-integral (relative to $\mathfrak{R}$}) if all its coefficients are in ${\mathfrak{R}}\otimes_{\mathbb{Z}}\tilde{\mathbb{Z}}_{p}$, where $\tilde{\mathbb{Z}}_{p}$ is the ring of rational numbers whose denominator is not divisable by $p$. Note that $\tilde{\mathbb{Z}}_{p}[\frac{1}{p}]=\mathbb{Q}$.
The following is an immediate consequence of the definition.
Proposition 1.9.
Consider a differential equation $y^{\prime}={\sf A}(z)y$ defined over $\mathfrak{R}$}. A solution of this differential equation with $y(z_{0})=y_{0},\ z_{0}\in{\mathfrak{R}},\ y_{0}\in{\mathfrak{R}}^{n}$, is $p$-integral if
$\frac{{\sf A}_{m}(z_{0})y_{0}}{m!}\in{\mathfrak{R}}\otimes_{\mathbb{Z}}\tilde{\mathbb{Z}}_{p},\ \ \forall m\in\mathbb{N}.$
In 1.3 we have proved that for an algebraic solution of a differential equation we have always $N\in{\mathfrak{R}}$ such that it is $p$-integral for all primes except a finite number and relative to the ring ${\mathfrak{R}}[\frac{1}{N}]$.
Definition 1.3.
We say that a differential equation $y^{\prime}={\sf A}(z)y$ defined over $\mathfrak{R}$} is $p$-integral (relative to $\mathfrak{R}$}) if
$\frac{\Delta(z)^{m}{\sf A}_{m}(z)}{m!}\in{\mathfrak{R}}\otimes_{\mathbb{Z}}\tilde{\mathbb{Z}}_{p}[z],\ \ \forall m\in\mathbb{N}.$
where $\Delta\in{\mathfrak{R}}[z]$ is a polynomial such the entries of $\Delta{\sf A}$ are also in ${\mathfrak{R}}[z]$.
If a differential equation is $p$-integral relative to $\mathfrak{R}$} then all its solutions at $z=z_{0}\in{\mathfrak{R}}$ with the initial value $y_{0}\in{\mathfrak{R}}^{n}$, are $p$-integral, relative to ${\mathfrak{R}}[\frac{1}{\Delta(z_{0})}]$, provided that $p$ is not invertible in this ring. However, the converse does not seem to be true, but the author is not aware of a counterexample. Note that the integer-valued polynomials like $\frac{z(z+1)}{2}$ have values in $\mathbb{Z}$ for all $z\in\mathbb{Z}$, however, they are not defined over $\mathbb{Z}$.
[…]
(1.14) $ny_{n}=\sum_{i+j=n-1}{\sf A}_{i}y_{j}.$
@@
We have also $ny(p)_{n}\equiv_{p}\sum_{i+j=n-1}{\sf A}_{i}y(p)_{j}$. For $n\leq p^{2}-1$ with $p\mid n$, the right hand side of (1.14) is divisible by $p$ and so $y_{n}$ can be computed in $\check{\mathfrak{R}}\otimes_{\mathbb{Z}}\tilde{\mathbb{Z}}_{p}$. However, for $n=p^{2}$, we observe that $y_{n}$ might have $p$ in its denominator and the argument breaks. One might try to give a recursive formula for the order of $p$ in the denominator of $y_{n}$. In case both $z_{0}$ and the differential equation are defined over $\mathbb{Q}$, the classical lower bound ${\rm ord}_{p}(y_{n})\geq\frac{-n}{p(p-1)}$ (see for instance [BostanCarusoRoques2024, Proposition 3.38]) does not see this phenomenon, for instance, it says that for $n=p^{2}-1$, the prime $p$ might divide the denominator of $y_{n}$, whereas by our recursion this does not happen.